# Energy-Efficient Transformer Inference: Optimization Strategies for Time Series Classification

### v2

@@ 7.4.1 Computational Resource Utilization
Comparative Analysis with State-of-the-Art Methods
@@
The experimental results reveal important insights regarding the trade-offs between model complexity, energy efficiency, and classification performance. Through careful analysis of these relationships, we observe that:
@@ Discussion
The results presented in this study highlight the trade-offs between model accuracy, inference time, energy consumption, and memory footprint when optimizing transformer models for time series classification. By systematically evaluating the effects of structured pruning and quantization across three datasets—RefrigerationDevices, ElectricDevices, and PLAID—we demonstrate that applying optimization strategies significantly enhances computational efficiency while maintaining satisfactory classification performance.
@@
Dataset-specific performance variations were also observed. The PLAID dataset exhibited the highest baseline accuracy (84.25% for T2), and even after optimization, models maintained high classification performance, with the lowest accuracy observed at 80.92% (after L2 pruning). In contrast, the RefrigerationDevices dataset, which had the lowest baseline accuracy (65.95% for T2), was more sensitive to optimization techniques, with accuracy dropping to 56.45% under L2 pruning. This suggests that the effectiveness of optimization strategies is dataset-dependent, and highly structured datasets like PLAID are more resilient to compression techniques.
The implications of these findings are significant for real-world deployments. Energy-efficient transformer models can substantially reduce operational costs and carbon footprints in applications such as predictive maintenance, appliance monitoring, and smart grid analytics. Additionally, in edge computing scenarios, where computational resources and power availability are limited, adopting a combination of quantization and pruning can enable real-time inference while maintaining acceptable accuracy levels.
@@ Conclusion
This study presents a systematic investigation of energy-efficient optimization techniques for transformer-based architectures in time series classification. Through extensive experimentation across three distinct datasets, we have demonstrated the effectiveness of combining hardware-aware optimization strategies while maintaining classification performance. Our experimental results reveal significant variations in model performance, with the PLAID dataset achieving 84.25% accuracy (T2 baseline) compared to 65.95% for RefrigerationDevices, highlighting the importance of dataset-specific optimization approaches.
The proposed framework demonstrates substantial improvements in energy efficiency, with static quantization achieving 29.14% energy savings while maintaining reasonable accuracy trade-offs. Notably, L1 pruning techniques achieved speed-ups of 1.63× compared to baseline models, while reducing energy consumption by 37.08%. The T2 architecture, despite higher computational complexity, provided superior accuracy-energy trade-offs across all datasets, suggesting the viability of optimizing larger models for energy-efficient deployment.
These findings have significant implications for deploying transformer models in resource-constrained environments. The variation in classification accuracy across datasets (60-84%) underscores the necessity of tailoring optimization strategies to specific application domains. Furthermore, the maintenance of high accuracy on the PLAID dataset (¿80%) post-optimization demonstrates the robustness of certain time series classification tasks to model compression, providing valuable insights for future implementations in edge computing scenarios.

### v3

@@ 5.3.1 Computational Resource Utilization
Comparative Analysis
@@
The experimental results reveal meaningful insights regarding the trade-offs between model complexity, energy efficiency, and classification performance. Through careful analysis of these relationships, we observe that:
@@ Discussion
Impact on Classification Accuracy
@@ Conclusion
Our study systematically investigates energy-efficient optimization strategies for transformer-based time series classification. By integrating structured pruning and quantization, we achieve significant improvements in inference speed and energy consumption while preserving robust classification performance. Notably, the proposed framework adapts to varying dataset characteristics, demonstrating that even models with higher computational complexity can be effectively compressed for deployment in resource-constrained environments. Our results underscore the need to tailor optimization strategies to specific application domains and offer valuable insights for developing scalable, sustainable deep learning solutions in edge computing scenarios.